# Multi-height analysis of rational points of toric varieties

### v1

@@ 2.1. Generalities about versal torsors
Originally, Colliot-Thélène and Sansuc defined universal torsors using a cohomological characterization (see [6, section 2]) that can be generalized over an noetherian base scheme $S$. For this paragraph, we essentially refer to [18, section 5].
@@
We suppose that $f:V\rightarrow S$ is a smooth proper surjective morphism of constant relative dimension with geometrically connected fibers such that $R^{2}f_{*}\mathcal{O}_{X}=0$ and the relative Picard functor $\operatorname{Pic}_{V/S}$ is étale locally the constant sheaf associated with an ordinary free abelian group of finite rank. We also suppose that $f$ has a rational section. We shall denote by $T_{\operatorname{NS}}$ the group scheme associated to $\operatorname{Pic}_{V/S}$ by duality. Then we have the following exact sequence, functorial in $S$ and $V$ (see [18, section 5, p.163] which generalize [6, 2.0.2]):
@@ A $T_{\operatorname{NS}}$-torsor $\mathcal{T}\rightarrow V$ is said to be universal if its class
$\chi([\mathcal{T}]):$ $H^{1}(V,\mathbf{G}_{m})\rightarrow H^{1}(V,G_{m})$
$[L^{\times}]\longmapsto[\mathcal{T}\overset{T_{\operatorname{NS}}}{\times_{\phi_{L}}}\mathbf{G}_{m}]$
where $\phi_{L}$ is the character of $T_{\operatorname{NS}}$ associated to $[L^{\times}]\in H^{1}(V,\mathbf{G}_{m})$ i.e $[L]\in\operatorname{Pic}(V)$.
$\mathcal{T}_{\mathbf{b}}=L_{1}^{\times}\times_{V}...\times_{V}L_{t}^{\times}$
The torus $T_{\operatorname{NS}}$ is isomorphic to $\mathbf{G}_{m}^{t}$ via the characters given by the basis $\mathbf{b}$. Hence we get a natural left action of $T_{\operatorname{NS}}$ over
$q:\mathcal{T}_{\mathbf{b}}\rightarrow V$
which makes $\mathcal{T}_{\mathbf{b}}$ a $T_{\operatorname{NS}}$-torsor over $V$.
In particular, provided that we have chosen the right convention for the contracted product, we find that $\chi([\mathcal{T}_{\mathbf{b}}])=\operatorname{Id}_{\operatorname{Pic}_{V/S}}$ (otherwise it will be $-\operatorname{Id}_{\operatorname{Pic}_{V/S}}$).
By [16, proposition 8.1], we get that the previous construction is independent, up to a $T_{\operatorname{NS}}$-equivariant isomorphism, to the choice of the basis of $\operatorname{Pic}(V)$ chosen.
Remark 2.2.
Let us recall the following proposition from [18, proposition 5.15], which is valid under more general assumptions that our hypothesis that $\operatorname{Pic}_{V/S}$ is a constant sheaf:
Proposition 2.3.
There is a universal torsor over $V$. Moreover the isomorphism classes of universal torsors over $V$ are parameterized by the $H^{1}(S,T_{\operatorname{NS}})$-orbits in $H^{1}(V,T_{\operatorname{NS}})$ defined by $\chi^{-1}(\operatorname{Id})$.
Suppose again that $S=\operatorname{Spec}(\mathbf{Z})$ and that $\operatorname{Pic}_{V/\mathbf{Z}}$ is the constant sheaf associated to a free abelian group of finite rank $t$. As $\mathbf{Z}$ is a principal ring, we have that $H^{1}(\mathbf{Z},T_{\operatorname{NS}})=\{1\}$, there is up to isomorphism a universal torsor $q:\mathcal{T}\rightarrow V$. Moreover, we have:
$q(\mathcal{T}(\mathbf{Z}))=V(\mathbf{Z}).$
2.2. Local heights defined over the universal torsors
We keep the hypothesis of the previous paragraph on $V$. We write $q:\mathcal{T}\rightarrow V$ for the universal torsor we have chosen (unique up to isomorphism by hypothesis). We fix $y_{0}\in\mathcal{T}(\mathbf{Q})$ and we denote by $x_{0}\in V(\mathbf{Q})$ its image via $q$.
We know ([17, remark 4.19]) that the universal torsor over $V_{\mathbf{Q}}$ is then characterized by a universal property as a pointed torsor.
Proposition 2.4.
For any pointed torsor $P\rightarrow V$ which send $p_{0}\in P(\mathbf{Q})$ to $x_{0}\in V(\mathbf{Q})$, under a multiplicative group $T$, there exists a unique group morphism $\varphi:T_{NS}\rightarrow T$ and a unique $\varphi$équivariant morphism $\Phi:\mathcal{T}\rightarrow P$ over $V$ which is compatible with the marqued point. That is to say the following diagram is commutative:
${{(\mathcal{T},y_{0})}}$${{(P,p_{0})}}$${{(V,x_{0})}}$$\scriptstyle{\exists!\Phi}$
We recall the construction of Peyre [17, construction 4.27] to define the local height associated to the universal torsor. Suppose we have a given heights system
$s:\operatorname{Pic}(V)\rightarrow\mathcal{H}(V)$
We fix a place $v_{0}\in M_{\mathbf{Q}}$. For any linde bundle $L$, we have by the previous proposition a unique morphism $\Phi_{L}:\mathcal{T}\rightarrow L^{\times}$ which is equivariant for the group morphism $\chi_{L}:T_{NS}\rightarrow\mathbf{G}_{m}$ given by $[L]\in X^{*}(T_{NS})=\operatorname{Pic}(V)$. This morphism is unique up to multiplication by a scalar $\lambda\in\mathbf{Q}^{\times}$.
$||-||_{L,v}:\mathcal{T}$ $(\mathbf{Q}_{v})\rightarrow\mathbf{R}$
$y\longmapsto\frac{||\phi_{L}(y)||_{v}}{||\phi_{L}(y_{0})||_{v}}\text{ si }v\neq v_{0}$
$y\longmapsto\frac{||\phi_{L}(y)||_{v_{0}}}{||\phi_{L}(y_{0})||_{v_{0}}}H(L,q(y_{0}))^{-1}$
Definition 2.5.
$H_{\mathcal{T},v}(y,L)=||y||_{L,v}^{-1}$
@@ 2.2. Local heights defined over the universal torsors
$H_{\mathcal{T},v}(y,L)=q_{v}^{-\langle h_{\mathcal{T},v}(y),L\rangle}.$
@@
Let $T$ be a torus over $\mathbf{Q}$, we denote $M=X^{*}(T)$ the $\operatorname{Gal}(\bar{\mathbf{Q}}/\mathbf{Q})$-lattice corresponding to the character group of $T$ and by $N=X_{*}(T)$ its dual for the usual pairing. For $v\in M_{\mathbf{Q}}$, we have a natural map:
@@
$|\chi(t)|_{v}=q_{v}^{\langle\log_{T,v}(t),\chi\rangle}.$
We can then write the following lemma ([17, lemme 4.30]):
@@ 2.3. Universal torsors for split, smooth and projective toric varieties
We keep the notations from Notations, let $X$ be a toric variety over $\mathbf{Q}$ (smooth, projective and split) and denote by $\Sigma$ an associated fan. In our particular case, the universal torsor $\mathcal{T}_{\Sigma}\rightarrow X$ can be described as an open subset of $\mathbf{A}^{\Sigma(1)}$ obtained by cutting out a certain closed subset $Z(\Sigma)$.
As in [7, Chapter 5, §1], we define for a cone of the fan $\sigma\in\Sigma$, the monomial:
@@
We shall give here an other description of the universal torsor $\mathcal{T}_{\Sigma}\rightarrow X$ that we shall not use in this paper but which is more intuitive. It is an open subset of $\mathbf{A}^{\Sigma(1)}$ obtained by cutting out a certain closed subset which can be described in term of primitive collections which are defined as follows (see [7, definition 5.1.5]):
@@
Hence $Z(\Sigma)$ can be described as a union of irreducible components as in [7, proposition 5.1.6]:
@@
Let $\beta:\operatorname{Div}_{T}(X)^{\vee}=\mathbf{Z}^{\Sigma(1)}\rightarrow N$ be the $\mathbf{Z}$-linear map which sends the vector of the canonical basis $e_{\rho}$ to $u_{\rho}$ the minimal generator of the edge $\rho$. Denote by $\Sigma_{0}$ the fan of $\mathcal{T}_{\Sigma}$ (see [7, proposition 5.19]). Both toric varieties $\mathcal{T}_{\Sigma}$ and $X$ are defined over $\mathbf{Z}$ via their fans. We state here the following essential proposition:
@@
The natural toric morphism $q:\mathcal{T}_{\Sigma}\rightarrow X$ over $\mathbf{Z}$ given by $\beta:(\mathbf{Z}^{\Sigma(1)},\Sigma_{0})\rightarrow(N,\Sigma)$ makes $\mathcal{T}_{\Sigma}$ the universal torsor (unique up to isomorphism) over $X$.
@@
$\chi_{D_{\rho}}:$ $T_{\Sigma(1)}\rightarrow\mathbf{G}_{m}$
@@
Note that we have $q^{-1}(T)=T_{\Sigma(1)}$.
Recall we have the following exact sequence associated to the toric variety $X$:
[…]
@@ 2.3.1. An alternative description of the system of heights over toric varieties
Using Kleiman’s criterion (see [9, theorem 1.27]), we can choose $(L_{1},..,L_{t})$ a basis of $\operatorname{Pic}(X)$ where each $L_{i}$ is an ample line bundle. Moreover by [7, theorem 6.1.16], because $X$ is smooth, $L_{i}$ ample implies $L_{i}$ very ample. We get the following lemma:
@@
We choose such basis. For $1\leqslant k\leqslant t$, we set $h^{0}(X,L_{k})=r_{k}$ and we take $\mathcal{M}_{k}\subset\operatorname{Div}_{T}(X)$ a family of $T$-invariant divisor such that:
@@
The fact that such family exists is a consequence of the description of the Cox Ring of the toric variety $X$. In fact, one can show that $\mathcal{M}_{k}=\pi^{-1}(L_{k})\cap\mathbf{N}^{\Sigma(1)}$.
@@
${\mathcal{T}_{\mathbf{b}}}$${\prod\limits_{i=1}^{t}\mathbf{A}^{r_{i}}\setminus\{0\}}$${X}$${\prod\limits_{i=1}^{t}\mathbf{P}^{r_{i}-1}}$
Let us denote by $A_{k}\in\mathcal{M}_{r_{k},m}(\mathbf{Z})$ the matrix whose lines are given by the coordinates of $D\in\mathcal{M}_{k}$ in the basis $\{D_{\rho}|\rho\in\Sigma(1)\}$.
@@
The $\mathbf{Z}$-linear map given by the matrix $\Bigg{[}\begin{smallmatrix}&A_{1}\\ &\vdots\\ &A_{t}\end{smallmatrix}\Bigg{]}$, that is to say:
induced the toric morphism of toric varieties:
which corresponds to the cartesian diagram 2.13. Moreover, the induced morphism by quotient
$X\rightarrow\prod\limits_{i=1}^{t}\mathbf{P}^{r_{i}-1}$
is again a toric morphism of toric varieties.
Hence using the height system previously defined for toric varieties, fixing the point $1=(1)_{\rho\in\Sigma(1)}$ of $\mathcal{T}_{\Sigma}(\mathbf{Q})$, we can describe the local height at the universal torsor $\mathcal{T}_{\Sigma}$. We use the functorial property 1.13 of the notion of heights system for toric morphism and we get that for any $v\in M_{\mathbf{Q}}$ and any $k\in\{1,..,t\}$ :
$H_{\mathcal{T},v}(y,L_{k})=\max\limits_{m\in\mathcal{M}_{k}}|X^{m}(y)|.$
In fact using the description of the universal torsor given by the cartesian diagram 2.13, we can deduce an other description of the $\mathbf{Z}$-points of $\mathcal{T}_{\Sigma}$:
Theorem 2.15.
$\mathcal{T}_{\Sigma}(\mathbf{Z})=\{y\in\mathbf{Z}^{\Sigma(1)}|\forall\ i\in\{1,..,t\},\ \gcd\limits_{D\in\mathcal{M}_{i}}(y^{D})=1\}$
@@
$H_{L_{i}}(q(y))=\max\limits_{D\in\mathcal{M}_{i}}(|y^{D}|)=H_{\mathcal{T},\infty}(y,L_{i}).$
As $L_{1},..,L_{t}$ form a basis of $\operatorname{Pic}(X)$, we get also the following corollary:
Corollary 2.16.
For any $y\in\mathcal{T}_{\Sigma}(\mathbf{Z})$, we have:
$h(q(y))=h_{\mathcal{T},\infty}(y).$

### v2

@@ 3.1. Generalities about universal torsors
Originally, Colliot-Thélène and Sansuc defined universal torsors using a spectral sequence (see [7, section 2]) that can be generalized over an noetherian base scheme $S$. For this paragraph, we essentially refer to [20, section 5].
@@
We suppose that $f:V\rightarrow S$ is a smooth proper surjective morphism of constant relative dimension with geometrically connected fibers such that $R^{2}f_{*}\mathcal{O}_{V}=0$ and the relative Picard functor $\operatorname{Pic}_{V/S}$ is étale locally the constant sheaf associated with an ordinary free abelian group of finite rank. We also suppose that $f$ has a rational section. We shall denote by $T_{\operatorname{NS}}$ the group scheme associated to $\operatorname{Pic}_{V/S}$ by duality. Then we have the following exact sequence, functorial in $S$ and $V$ (see [20, section 5, p.163] which generalizes [7, 2.0.2]):
$0\rightarrow H^{1}(S,T_{\operatorname{NS}})\xrightarrow{i_{1}}H^{1}(V,T_{\operatorname{NS}})\xrightarrow{\chi}\operatorname{Hom}_{\operatorname{grp}}(\operatorname{Pic}_{V/S},\operatorname{Pic}_{V/S})\xrightarrow{\partial}H^{2}(S,T_{\operatorname{NS}})$
Definition 3.1.
$\mathcal{T}]\in H^{1}(V,T_{\operatorname{NS}})$
is sent to $\operatorname{Id}_{\operatorname{Pic}_{V/S}}$ via $\chi$.
In the case we will be interested, $S=\operatorname{Spec}(\mathbf{Z})$, and our scheme will be split that is to say $\operatorname{Pic}_{V/S}$ is the constant sheaf associated with a free abelian group of finite rank. With this hypothesis, we have that for any $S$-scheme $T$, $\operatorname{Pic}_{V/S}(T)=H^{1}(V\underset{S}{\times}T,\mathbf{G}_{m})$ and $\operatorname{Hom}_{\operatorname{grp}}(\operatorname{Pic}_{V/S},\operatorname{Pic}_{V/S})=\operatorname{Hom}_{\operatorname{grp}}(\operatorname{Pic}(V),\operatorname{Pic}(V))$. Moreover, the group morphism $\chi$ can be defined for any $\mathcal{T}]\in H^{1}(X,T_{\operatorname{NS}})$ as follows (see [21, Definition 2.3.2 and Theorem 2.3.6]):
$\chi([\mathcal{T}]):\$$H^{1}(V,\mathbf{G}_{m})\rightarrow H^{1}(V,\mathbf{G}_{m})$
$[L^{\times}]\longmapsto[\mathcal{T}\overset{T_{\operatorname{NS}}}{\times_{\chi_{L}}}\mathbf{G}_{m}]$
where $\chi_{L}$ is the character of $T_{\operatorname{NS}}$ associated to $L^{\times}]\in H^{1}(V,\mathbf{G}_{m})$ i.e. $L]\in\operatorname{Pic}(V)$.
$\mathcal{T}_{\mathbf{b}}=L_{1}^{\times}\times_{V}...\times_{V}L_{t}^{\times}$
where an element $t\in T_{\operatorname{NS}}$ acts on $L_{i}^{\times}$ by multiplication by $L_{i}](t)\in\mathbf{G}_{m}$. The torus $T_{\operatorname{NS}}$ is isomorphic to $\mathbf{G}_{m}^{t}$ via the characters given by the basis $\mathbf{b}$. Hence we get a natural left action of $T_{\operatorname{NS}}$ over
$q:\mathcal{T}_{\mathbf{b}}\rightarrow V$
which makes $\mathcal{T}_{\mathbf{b}}$ a $T_{\operatorname{NS}}$-torsor over $V$. By [18, proposition 8.1], we get that the previous construction is independent of the choice of the basis of $\operatorname{Pic}(V)$ because it is unique up to $T_{\operatorname{NS}}$-equivariant isomorphism. To simplify the notations, we shall denote it $\mathcal{T}\rightarrow V$. If each $L_{i}$ has a non constant global section, we can show that $\chi([\mathcal{T}])=-\operatorname{Id}_{\operatorname{Pic}_{V/S}}$.
Remark 3.2.
Let us recall the following proposition from [20, proposition 5.15], which is valid under our assumptions that $\operatorname{Pic}_{V/S}$ is a constant sheaf and $V$ has a rational point:
Proposition 3.3.
There is a universal torsor over $V$. Moreover the isomorphism classes of universal torsors over $V$ are parametrized by the $H^{1}(S,T_{\operatorname{NS}})$-orbits in $H^{1}(V,T_{\operatorname{NS}})$ defined by $\chi^{-1}(\operatorname{Id}_{\operatorname{Pic}_{V/S}})$.
Suppose again that $S=\operatorname{Spec}(\mathbf{Z})$ and that $\operatorname{Pic}_{V/\mathbf{Z}}$ is the constant sheaf associated to a free abelian group of finite rank $t$. As $\mathbf{Z}$ is a principal ring, we have that $H^{1}(\mathbf{Z},T_{\operatorname{NS}})=\{1\}$. Thus there is up to isomorphism a unique universal torsor $q:\mathcal{T}\rightarrow V$. Moreover, we have:
$q(\mathcal{T}(\mathbf{Z}))=V(\mathbf{Z}).$
3.2. Local heights defined over the universal torsors
We keep the hypotheses of the previous paragraph on $V$. By [19, remark 4.19], the universal torsor over $V_{\mathbf{Q}}$ is characterized by a universal property as a pointed torsor.
Definition 3.4.
A pointed variety over $\mathbf{K}$ is a variety $V$ equipped with a chosen rational point $x\in V(\mathbf{K})$. A pointed torsor over $V$ is a torsor $P\xrightarrow{\pi}V$ with a rational point $t\in P(\mathbf{K})$ such that $\pi(t)=x$.
We write $q:\mathcal{T}\rightarrow V$ for the universal torsor we have chosen. We fix $y_{0}\in\mathcal{T}(\mathbf{Q})$ and we denote by $x_{0}\in V(\mathbf{Q})$ its image via $q$. We can now state the following proposition:
Proposition 3.5.
For any pointed torsor $P\rightarrow V$ which sends $p_{0}\in P(\mathbf{Q})$ to $x_{0}\in V(\mathbf{Q})$, under a multiplicative group scheme $T$, there exists a unique group morphism $\varphi:T_{NS}\rightarrow T$ and a unique $\varphi$ -équivariant morphism $\Phi_{P}:\mathcal{T}\rightarrow P$ over $V$ which is compatible with the marked point. That is to say the following diagram is commutative:
$(\mathcal{T},y_{0})$$(P,p_{0})$$(V,x_{0})$$\scriptstyle{\exists!\Phi_{P}}$
We recall the construction of Peyre [19, construction 4.27] to define the local height associated to the universal torsor. Suppose we have a given system of heights
$s:\operatorname{Pic}(V)\rightarrow\mathcal{H}(V).$
We fix a place $v_{0}\in M_{\mathbf{Q}}$. For any line bundle $L$, we have by the previous proposition a unique morphism $\Phi_{L}:\mathcal{T}\rightarrow L^{\times}$ which is equivariant for the group morphism $\chi_{L}:T_{NS}\rightarrow\mathbf{G}_{m}$ given by $L]\in X^{*}(T_{NS})=\operatorname{Pic}(V)$. This morphism is unique up to multiplication by a scalar $\lambda\in\mathbf{Q}^{\times}$.
$||-||_{L,v}:\mathcal{T}$ $(\mathbf{Q}_{v})\rightarrow\mathbf{R}$
[…]
@@ 3.2. Local heights defined over the universal torsors
$H_{\mathcal{T},v}(y,L)=q_{v}^{\langle h_{\mathcal{T},v}(y),L\rangle}.$
@@
Let $T$ be a split torus over $\mathbf{Q}$, we denote $M=X^{*}(T)$ the lattice corresponding to the character group of $T$ and by $N=X_{*}(T)$ its dual for the usual pairing. For $v\in M_{\mathbf{Q}}$, we have a natural map:
@@
$\langle\log_{T,v}(t),\chi\rangle:=-\log_{q_{v}}\left|\chi(t)\right|_{v}.$
With our convention, we can rewrite the following lemma ([19, lemma 4.30]):
@@ 3.3. Universal torsors for split, smooth and projective toric varieties
We keep the notations from Notations. Let $X$ be a toric variety over $\mathbf{Q}$ (smooth, proper and split) and denote by $\Sigma$ an associated fan. In our particular case, the universal torsor $\mathcal{T}_{\Sigma}\rightarrow X$ can be described as an open subset of $\mathbf{A}^{\Sigma(1)}$ obtained by cutting out a certain closed subset $Z(\Sigma)$.
As in [8, chapter 5, §1], we define for a cone of the fan $\sigma\in\Sigma$, the monomial:
@@
Remark 3.10.
There is a more intuitive description of the universal torsor $\mathcal{T}_{\Sigma}\rightarrow X$ as an open subset of $\mathbf{A}^{\Sigma(1)}$ obtained by cutting out a certain closed subset which can be described in term of primitive collections which are defined as follows (see [8, definition 5.1.5]):
@@
Hence $Z(\Sigma)$ can be described as a union of irreducible components as in [8, proposition 5.1.6]:
@@
Let $\beta:\operatorname{Div}_{T}(X)^{\vee}=\mathbf{Z}^{\Sigma(1)}\rightarrow N$ be the $\mathbf{Z}$-linear map which sends the vector of the canonical basis $e_{\rho}$ to $u_{\rho}$ the minimal generator of the edge $\rho$. Denote by $\Sigma_{0}$ the fan of $\mathcal{T}_{\Sigma}$ (see [8, proposition 5.19]). Both toric varieties $\mathcal{T}_{\Sigma}$ and $X$ are defined over $\mathbf{Z}$ via their fans. We state here the following essential proposition:
@@
The natural quotient morphism $q:\mathcal{T}_{\Sigma}\rightarrow X$ over $\mathbf{Z}$ where $T_{\operatorname{NS}}$ acts on the coordinate associated to $\rho\in\Sigma(1)$ by multiplication by $D_{\rho}](t)^{-1}$ makes $\mathcal{T}_{\Sigma}$ the universal torsor (unique up to isomorphism) over $X$.
@@
$\chi_{D_{\rho}}:\$$T_{\Sigma(1)}\rightarrow\mathbf{G}_{m}$
@@
Note that we have $q^{-1}(T)=T_{\Sigma(1)}$. Recall we have the following exact sequence associated to the toric variety $X$:
$0\rightarrow M\xrightarrow{\beta^{*}}\operatorname{Div}_{T}(X)\xrightarrow{\pi}\operatorname{Pic}(X)\rightarrow 0.$
@@
Proof.
Using Kleiman’s criterion (see [10, theorem 1.27]), we can choose $(L_{1},..,L_{t})$ a basis of $\operatorname{Pic}(X)$ where each $L_{i}$ is an ample line bundle. Moreover by [8, theorem 6.1.16], because $X$ is smooth, $L_{i}$ ample implies $L_{i}$ very ample. We have proven the lemma. ∎
We choose such basis. For $1\leqslant k\leqslant t$, we set $h^{0}(X,L_{k})=r_{k}$ and we take $\mathcal{M}_{k}=\pi^{-1}(L_{k})\cap\mathbf{N}^{\Sigma(1)}\subset\operatorname{Div}_{T}(X)$. Thanks to the description of the Cox Ring of the toric variety $X$, it is a family of $T$-invariant divisor such that:
@@
[…]
where $\mathbf{b}=(L_{1},..,L_{t})$ and the horizontal maps are toric morphism.
Proof.
Let us denote by $A_{k}\in M_{r_{k},m}(\mathbf{Z})$ the matrix whose lines are given by the coordinates of $D\in\mathcal{M}_{k}$ in the basis $\{D_{\rho}|\rho\in\Sigma(1)\}$. The $\mathbf{Z}$-linear map given by the matrix $\Bigg[\begin{smallmatrix}&A_{1}\\ &\vdots\\ &A_{t}\end{smallmatrix}\Bigg]$, that is to say:
induces the toric morphism of toric varieties over $\mathbf{Z}$:
which corresponds to the cartesian diagram 3.15. Moreover, the induced morphism by quotient
$X\rightarrow\prod\limits_{i=1}^{t}\mathbf{P}^{r_{i}-1}$
is again a toric morphism of toric varieties. ∎
Recall that we have equipped the toric variety $X$ with its natural system of heights as in convention 2.12. We fix the point $1=(1)_{\rho\in\Sigma(1)}$ of $\mathcal{T}(\mathbf{Q})$ and $q(1)=1\in X(\mathbf{Q})$. In this paragraph, we shall explain how to lift the height of a rational point $P\in X(\mathbf{Q})$ to the universal torsor, that is to say, we shall prove the following theorem:
Theorem 3.16.
Let $y\in\mathcal{T}(\mathbf{Z})$ and $P=q(y)\in X(\mathbf{Q})$. We have:
$h(P)=h_{\mathcal{T},\infty}(y).$
@@ 3.3.1. An alternative description of the system of heights over toric varieties
The system of heights $s_{X}:\operatorname{Pic}(X)\rightarrow\mathcal{H}(X)$ used in [20] on a toric variety $X$ is functorial for the morphisms of split toric varieties.
That is to say, for $X\xrightarrow{f}Y$ a toric morphism of split toric varieties (i.e. compatible with the actions of the tori), the following diagram is commutative:
Proof.
For a split toric variety $X$, we denote by $T_{\operatorname{NS},X}$ its Néron–Severi torus and by $\mathcal{T}_{X}$ its universal torsor, as defined previously. According to [13, section 3], a toric morphism
$X\xrightarrow{f}Y$
between toric varieties lifts to a morphism of toric varieties
$\mathcal{T}_{X}\xrightarrow{\tilde{f}}\mathcal{T}_{Y}$
which is equivariant with respect to a morphism of tori $T_{\operatorname{NS},X}\rightarrow T_{\operatorname{NS},Y}$.
In [20, theorem 9.5], Salberger constructs a canonical section
$X(K_{v})\xrightarrow{\check{\psi}_{X}}\mathcal{T}_{X}(K_{v})/T_{\operatorname{NS},X}(K_{v})_{\mathrm{cp}}$
that, by construction, behaves functorially in $X$ with respect to toric morphisms. Here, $T_{\operatorname{NS},X}(K_{v})_{\mathrm{cp}}$ denotes the maximal compact subgroup of $T_{\operatorname{NS},X}(K_{v})$.
The proposition then follows from the fact that $T(K_{v})$ is dense in $X(K_{v})$ and that, according to [20, proposition 9.7], we have the equality
$1(P)||_{D}=|\check{\psi}(P)|_{D},$
where the definition of $|-|_{D}$ is given in [20, notations 9.6]. Indeed, for a divisor $D$ on $Y$, we have
[…]
We can describe the local height at the universal torsor $\mathcal{T}_{\Sigma}$. We use the functorial property 3.17 of the notion of system of heights for toric morphism and we get that for any $v\in M_{\mathbf{Q}}$ and any $k\in\{1,..,t\}$ :
$H_{\mathcal{T},v}(y,L_{k})=\max\limits_{D\in\mathcal{M}_{k}}|X^{D}(y)|_{v}.$
In fact using the description of the universal torsor given by the cartesian diagram 3.15, we can deduce an other description of the $\mathbf{Z}$-points of $\mathcal{T}_{\Sigma}$:
Proposition 3.18.
$\mathcal{T}_{\Sigma}(\mathbf{Z})=\{y\in\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}_{\Sigma}(\mathbf{Q})|\forall\ i\in\{1,..,t\},\ \gcd\limits_{D\in\mathcal{M}_{i}}(X^{D}(y))=1\}.$
@@
$H_{L_{i}}(q(y))=\max\limits_{D\in\mathcal{M}_{i}}(|X^{D}(y)|)=H_{\mathcal{T},\infty}(y,L_{i}).$
As $L_{1},..,L_{t}$ form a basis of $\operatorname{Pic}(X)$, we have proven theorem 3.16.